OpenStudy (albert0898):

Please explain! What is the multiplicative inverse of 3i?

OpenStudy (xapproachesinfinity):

what do you think

OpenStudy (xapproachesinfinity):

oh by the way is that \[3i \]

OpenStudy (albert0898):

Yes, it's 3i

OpenStudy (xapproachesinfinity):

recall that if 1/a is a multiplicative inverse of a then \[a\times\frac{1}{a}=1\]

OpenStudy (anonymous):

3/i

OpenStudy (xapproachesinfinity):

so can you find such number that satisfies this condition

OpenStudy (albert0898):

Answer is actually -i/3, but how do I figure that out?

OpenStudy (anonymous):

because technically its 3 times i so its 3i, so the multiplicative inverse of 3i is division, so its 3 over i

OpenStudy (xapproachesinfinity):

i would agree with that answer -i/3 see the condition i wrote and reflect on it

OpenStudy (albert0898):

@Saisuke<3 i stands for imaginary, it's not just a variable

OpenStudy (xapproachesinfinity):

thing of 3i as a 1/a is the inverse of a then we should have 1/3i yes?

OpenStudy (xapproachesinfinity):

think*

OpenStudy (albert0898):

But that wouldn't make any sense as the answer is -i/3

OpenStudy (xapproachesinfinity):

well it makes perfect sense actually if 1/3i is the inverse of 3i you just need to do some rationalizing the denominator to get rid of i

OpenStudy (mathmate):

@Albert0898 -i/3 and 1/3i are the same number, like @xapproachesinfinity said. We prefer the first because we don't like i in the denominator, just like we write \(\dfrac{\sqrt 2}{2}\) instead of \(\dfrac{1}{\sqrt 2}\) So by multiplying 1/3i by i top and bottom, we get i/(3(-1)=-i/3

OpenStudy (albert0898):

Ah, now I get it. I knew how to the inverse and I got 1/3i and that wasn't a choice. Thank you all!

OpenStudy (xapproachesinfinity):

np!